Question 28

Calculate the price change for a 1-percent decrease in market yield for the following bond: par = $1,000; coupon rate = 7 percent, paid semi-annually; market yield = 7 percent; term to maturity = 9 years. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 4 decimal places, e.g. 1,564.2556.)

Change in price $

Practice Question 7

Which bond is most likely to see the smallest fluctuations in its market price when there is a change in market yields?

A bond that is far from its maturity date.

A long term bond that was just issued.

Time to maturity has no effect on the size of a bond’s market price fluctuations when the market yields changes.

A bond that is close to its maturity date.

Question 36

The price of a ten-year semi-annual pay bond with a par value of $1,000 and a 7.5 percent annual coupon and yield to maturity of 8.25 percent is closest to:

$949.60

$950.24

$1,051.48

$1,052.11
Answer #1

28. Semi annual Coupon =7%*1000/2 =35
Semi annual YTM =7%/2 =3.50%
Number of Periods =9*2 =18
Price when coupon rate and YTM are same then Price =1000

Price with new YTM of 6%
Semi annual YTM =6%/2 =3%
Price of Bond =PV of Coupons+PV of Par Value =35*((1-(1+3%)^-18)/3%)+1000/(1+3%)^18 =1068.77
Price change =1068.77-1000 =68.77

29. Option d is correct option.
A bond that is close to its maturity date.

30. Semi annual Coupon =7.5%*1000/2 =37.50
Semi annual YTM =8.25%/2 =4.125%
Number of Periods =10*2 =20
Price of Bond =PV of Coupons+PV of Par Value =37.50*((1-(1+4.125%)^-20)/4.125%)+1000/(1+4.125%)^20 =949.60
(Option a is correct option)